# System for Systematic Literature Review Using Multiple AI agents: Concept and an Empirical Evaluation

### v1

@@ Abstract
Systematic Literature Reviews (SLRs) have become the foundation of evidence-based studies, enabling researchers to identify, classify, and combine existing studies based on specific research questions. Conducting an SLR is largely a manual process. Over the previous years, researchers have made significant progress in automating certain phases of the SLR process, aiming to reduce the effort and time needed to carry out high-quality SLRs. However, there is still a lack of AI agent based models that automate the entire SLR process. To this end, we introduce a novel multi-AI agent model designed to fully automate the process of conducting an SLR. By utilizing the capabilities of Large Language Models (LLMs), our proposed model streamlines the review process, enhancing efficiency and accuracy. The model operates through a user-friendly interface where researchers input their topic, and in response, the model generates a search string used to retrieve relevant academic papers. Subsequently, an inclusive and exclusive filtering process is applied, focusing on titles relevant to the specific research area. The model then autonomously summarizes the abstracts of these papers, retaining only those directly related to the field of study. In the final phase, the model conducts a thorough analysis of the selected papers in relation to predefined research questions. This paper details the development of the model, and its operational framework, and demonstrates how it significantly reduces the time and effort traditionally required for SLR while ensuring a high level of comprehensiveness and precision. We also conducted an evaluation of the proposed model by sharing it with ten competent software engineering researchers for testing and analysis. The researchers expressed strong satisfaction with the proposed model and provided feedback for further improvement. In the future, we plan to engage 50 practitioners and researchers to evaluate our model. Additionally, we aim to present our model to the audience at the SANER 2024 conference in Rovaniemi (Finland) for further testing, analysis, and feedback collection. The code for this project can be found on the GitHub repository at https://github.com/GPT-Laboratory/SLR-automation.
@@ 1 Introduction
In this context, our proposed model utilizes the capabilities of LLMs to automate the whole SLR process. We developed a multi-AI agent model that automates each step of the SLR, from the initial literature search to the final analysis. The model begins with a simple user input–researchers enter their topic into a designated text box. This input is then processed by the LLM, which generates a precise search string tailored to retrieve the most relevant academic papers. The model’s next phase involves an intelligent filtering mechanism. It applies an inclusive and exclusive theory, screening titles and abstracts to retain only those studies that are directly relevant to the specified research area.
The final stage of our model autonomously summarizes the abstracts of the selected papers, ensuring that only content pertinent to the research questions is retained. It introduces a level of precision and consistency in data analysis that is challenging to achieve manually. Finally, the model conducts an in-depth analysis of the selected papers, aligning its findings directly with the research questions. This comprehensive approach ensures that the final output is not only a reflection of the vast array of literature available but also a focused and relevant resource tailored to the specific needs of the researcher. Our model, therefore, stands as a testament to the potential of integrating advanced AI technologies in academic research methodologies.
We also evaluated the efficiency and accuracy of our proposed model by sharing it with ten proficient software engineering researchers for a comprehensive test and analysis. The feedback received was overwhelmingly positive, highlighting the model’s effectiveness and paving the way for further enhancements. Looking ahead, we aim to expand our evaluation by involving 50 additional practitioners and researchers. Furthermore, we intend to showcase our model at the upcoming SANER 2024 conference in Rovaniemi (Finland), seeking to broaden its testing and analysis while gathering valuable feedback from a wider audience. This step is crucial in refining our model and ensuring its applicability and robustness in diverse real-world scenarios.
Our contribution can be summarized as follows:
  • We propose a novel multi-AI agent model that utilizes LLMs to automate the SLR process, significantly enhancing efficiency and accuracy.
  • Our model was evaluated by ten experienced software engineering researchers and practitioners, confirming its effectiveness and gathering insights for further refinement.
  • We plan to extend the evaluation to 30 additional practitioners and researchers and present the model at the SANER 2024 conference in Rovaniemi (Finland) for wider testing and feedback.
@@ 2 Related Work
Despite these advancements in automating the SLR process, there remains a notable gap in the complete automation of SLRs using LLMs. Addressing this gap, we have developed a novel approach: a multi-agent model based on LLMs. This innovative model is designed to fully automate the SLR process, utilizing the advanced capabilities of LLMs to efficiently manage and synthesize vast amounts of data, which is a significant step forward in the field of automated literature reviews.
@@ 3 Research Method
This research aims to investigate how an LLM-based multi-agent model can be utilized to automate the entire process of SLRs. We also outline the methodology for testing and analyzing the capabilities of the proposed model. Below, we discuss how our LLM-based multi-agent model collaborates and performs such tasks. We have formulated the following research questions (RQs):
RQ1. How does a LLM-based multi-agent system transform traditional methodologies to automate the systematic literature review process in SE?
Motivation: The motivation for the research question arises from the need to enhance the efficiency and effectiveness of literature review processes in the rapidly evolving field of SE. Traditional methods of conducting literature reviews are often time-consuming and labor-intensive, potentially leading to delays in research progress and the dissemination of new knowledge. The integration of LLMs promises a paradigm shift, potentially automating and streamlining these processes. By exploring the transformation brought about by an LLM-based multi-agent system, this research seeks to reduce the time and effort required for comprehensive literature reviews and also to increase the accuracy and scope of these reviews. This could result in more timely and informed research outcomes in SE, a field where staying abreast of current trends, methodologies, and discoveries are crucial for technological advancement and innovation.
RQ2. How can the efficiency and accuracy of the proposed LLM-based multi-agent model be evaluated?
Motivation: The motivation behind this research question is based on the critical need to validate and quantify the performance of the proposed model, specifically in conducting SLRs. With the introduction of sophisticated models like LLM-based multi-agent systems, establishing rigorous evaluation criteria to assess their real-world applicability and reliability becomes imperative. This question addresses the necessity of a proposed model that can systematically measure its efficiency and accuracy in selecting and interpreting relevant literature. Evaluating the proposed model is crucial to ensure that integrating such models into academic workflows enhances, rather than compromises, the quality of research outputs.
3.1 LLM-Based Assisted Systematic Literature Review
The section focuses on research methodology for developing an LLM-based multi-agent model. This model is specifically engineered to automate the entire process of conducting SLRs. The innovation lies in its ability to transform a given research topic into a comprehensive review through a series of automated, interconnected steps. Each step is managed by a specialized agent within the model, working collaboratively to ensure a seamless and efficient literature review process. In Figure LABEL:fig:_multi-agent_system, we illustrate how agents collaborate with each other to generate a response. Below, we also provide the detailed functionality of each agent within this multi-agent system.
@@ 3.1.1 Planner agent
The first agent in our model is dedicated to generating research questions and purpose and search strings. Upon receiving a research objective from the end-user, For the search string we provide research questions and objective to the agent, it generates a search string. this agent employs advanced language understanding algorithms to interpret the topic’s key elements. The LLM-based algorithm, designed for deep semantic understanding, analyzes the topic to extract key concepts, themes, and terminologies. It then utilizes its extensive training on diverse textual data to construct a precise and comprehensive search string. This string is formulated by combining relevant keywords, synonyms, and technical terms that capture the essence of the research question. Furthermore, the algorithm is adept at understanding context and varying semantic structures, enabling it to refine the search string to match specific research domains. The generated search string is crucial in accurately retrieving relevant literature from various academic databases. By ensuring that the initial search is both thorough and focused, the agent significantly enhances the efficiency and quality of the literature collection process. This sets a solid foundation for the subsequent stages of the SLR, where the depth and breadth of the collected literature play a crucial role.
@@ 3.1.3 Data extraction agent
The third agent in our model is tasked with refining the literature using inclusion and exclusion criteria based on the research objectives. Initially, it employs our LLM algorithm to analyze the titles of retrieved papers, discerning their relevance to the research topic. This step involves text analysis, where the LLM algorithm identifies key terms and concepts that align with the research objectives. By applying these predefined rules, the agent effectively filters out irrelevant material, ensuring the literature review remains focused and pertinent to the research questions.
@@ 3.1.4 Data compilation agent
The final agent in our multi-agent model is responsible for analyzing the synthesized data in relation to the research questions and objectives. It assesses trends, identifies gaps in the literature, and draws conclusions based on the aggregated information. This agent also prepares a report that summarizes the findings of the literature review, providing a clear and concise overview of the research landscape for the given topic.
Each agent in the LLM-based multi-agent model plays a vital role in automating the systematic literature review process. From generating search strings to reporting findings, the agents work in a coordinated manner to ensure a thorough, efficient, and accurate review. This methodology represents a significant advancement in the way SLRs are conducted, offering a more streamlined and effective approach to academic research.
3.2 Performance Validation
In this project, we engaged the expertise of 10 researchers to evaluate the efficiency and performance of our proposed model. These professionals were approached from various industries and research groups to ensure a diverse range of perspectives. Our evaluation methodology was anchored in a practice-oriented framework, focusing on rigorous scrutiny of the model’s utility and efficacy by seasoned industry professionals. This approach guaranteed a detailed and perceptive evaluation, factoring in the real-world applications and needs of experts in the corresponding fields. The inclusion of insights from a diverse group of experienced contributors was instrumental in our aim to deliver a comprehensive overview of the model’s functionality and its possible influence in the industry.
3.2.1 Professional-based evaluation
In the validation stage of our research, we collaborated with ten experts from diverse sectors, including academia and industry, to evaluate the efficacy of our LLM-based model. This selection was strategic, and aimed at capturing a wide range of perspectives and insights on the model’s performance across various professional contexts. To facilitate a thorough evaluation, each expert was provided with access to the model, accompanied by comprehensive instructions on its intended use and capabilities. This approach ensured that they were adequately equipped to conduct an in-depth assessment of the model within their respective domains.
Participant’s selection: Initially, we conducted a search for suitable researchers from various research groups and expanded our outreach to individuals through social networks. We reached out to researchers via their ResearchGate and Google Scholar profiles. From these platforms, we recruited five participants, and the remaining five were identified through social networking sites and professional contacts. Consequently, we assembled a group of ten participants, referred to as P1 to P10 in Table 1, for the evaluation of our proposed model. Table 1 demonstrates that these participants come from a variety of fields, such as Software Engineering, and Machine Learning/Deep Learning Development.
Data collection: The methodology employed a systematic approach to gather feedback data. Participants were tasked with integrating the model into their research and providing a topic relevant to their expertise. They then compared the results generated by our model with those obtained manually. To collect their feedback, we implemented a comprehensive feedback mechanism that enabled them to document their experiences, observations, and critiques systematically. This mechanism was geared towards obtaining detailed responses about the model’s efficiency, user-friendliness, and accuracy in data analysis. The format of the feedback was meticulously crafted to extract in-depth qualitative insights, thereby facilitating the assessment of quantifiable elements of the model’s performance.
Data analysis: In the final phase of our validation process, analyzing participant feedback was essential for ongoing model improvement. Participants were allowed to choose any data source for input. We gathered and scrutinized their feedback to understand how the model performed in various scenarios. To evaluate the model’s performance, we used an extensive Likert scale ranging from ’Not Satisfied’ to ’Excellent,’ with intermediate options like ’Fair,’ ’Satisfactory,’ ’Good,’ and ’Very Good.’ This scale provided a nuanced spectrum for evaluation, enabling precise and comprehensive grading of the model’s effectiveness. Our iterative approach was driven by a focus on enhancing the model’s functionality, user experience, and overall efficacy, aiming to fulfill the varied needs and expectations of professionals in qualitative data analysis.
@@ 4 Results
This section presents the results obtained from implementing an AI-agent-based model aimed at automating the SLR process in SE. The findings are detailed in accordance with two primary research questions (RQs) that steered the creation and evaluation of the model.
4.1 LLM Based Multi-Agent Model (RQ1)
Our research introduces an LLM-based multi-agent model that redefines the conventional approach to SLRs in SE. The multi-agent model developed for automating the SLR process has demonstrated its efficacy through a structured and sequential workflow. The process begins with the input of a research topic, as depicted in Figure 2.
@@ 4.1 LLM Based Multi-Agent Model (RQ1)
Upon receiving the topic, the model systematically generates a pertinent set of research questions. As illustrated in Figure 2, it formulated questions like ’How have large language models been utilized in various aspects of the software development process?’ and ’What challenges and limitations exist in the adoption and implementation of large language models in software development?’ These questions play a crucial role in guiding the literature search and analysis process. Subsequent to the research question formulation, the model proceeds to generate a search string. In this case, the search string “large language models OR software development” was created, coupled with a specified year to narrow down the search scope, thereby enhancing the relevance and precision of the search results.
@@
The subsequent phase involves retrieving papers that match the generated search string. This feature of the model is specifically designed to fetch research papers from various databases, as demonstrated by the list of papers shown in Figure 3. For this demonstration, we focused solely on papers published in the year 2023, setting the model to retrieve only 10 papers from that year, all relevant to this field. The tool efficiently compiles relevant information such as the title, author, publication URL, journal name, DOI, paper type, affiliation country, and the affiliation institution. Moreover, the model is equipped with the capability to apply inclusive and exclusive criteria based on titles, which further refines the search results to ensure only the most pertinent literature is considered for review. As shown in Figure 4, only three papers were selected for an in-depth analysis.
Finally, the model extracts data based on the formulated RQs. This advanced feature is exemplified in the demo where detailed answers are provided for the previously generated RQs. For instance, the answer to the first RQ discusses the varied applications of large language models in the software development lifecycle and highlights specific instances of their use, like the inference from the paper “InferLink End-to-End Program Repair with Large Language Models.”
@@
In conclusion, the automated SLR tool has showcased its ability to streamline the laborious process of literature review, from defining the scope of the research to extracting and synthesizing data pertinent to the research questions. The demonstration affirms the model’s potential in significantly reducing the time and effort conventionally required for conducting systematic literature reviews.
@@
4.2 Evaluation Result (RQ2)
The empirical evaluation of our tool was conducted by involving ten researchers and practitioners from diverse backgrounds within the SE community. Their engagement with the model provided a comprehensive perspective on its practical utility and user experience. The feedback was overwhelmingly positive, with 80% of the participants approving the tool’s functionality and recognizing its contribution to simplifying the SLR process.
Despite the general consensus on the model’s efficacy, 20% of the participants recommended improvements. Specific suggestions highlighted the need for a more nuanced interpretation of complex research queries and the generation of more refined search strings. This constructive feedback is invaluable as it directs the focus toward enhancing the model’s interpretative algorithms and its ability to handle ambiguous or multifaceted research questions.
In pursuit of continuous improvement, the model is scheduled for further exposure and evaluation. The SANER 2024 conference in Rovaniemi (Finland) presents an opportunity for a wide array of feedback from the SE research community, which will be instrumental in the model’s iterative development. Furthermore, a large-scale testing initiative is planned, where the model will be disseminated to a group of 50 researchers and practitioners for extensive evaluation.
This forthcoming phase is expected to yield deeper insights into the model’s generalizability and performance across various domains within SE. It will also help to pinpoint any subtle nuances in SLR processes that the model needs to accommodate. The comprehensive feedback will be integral to refining the model, ensuring that the final version is not only effective and efficient but also versatile and user-friendly. The ultimate objective is to deliver a robust, universally applicable tool that standardizes and automates SLRs, contributing to the advancement of research methodologies in SE.
Table 1: Researcher’s demography and their assessment
ID Role of Practitioner Experience Performance Feedback Suggestion
P1
AI Researcher
5 Years Very Good
Suggested enhancements in the number of years for paper selection.
Include more steps for refinement.
P2
SE Researcher
8 Years Excellent
Impressed with accuracy, recommended UI improvements.
User interface could be more intuitive.
P3
Empirical Researcher
10 Years Good
Advised on incorporating general terminology support.
Include more steps for data analyze.
P4
Theoretical Researcher
3 Years Satisfactory
Suggested to improve search string step.
Improve the paper selection criteria.
P5
Data Scientist in SE
6 Years Very Good
Pleased with model, suggested speed optimizations.
Increase processing speed.
P6
SE Researcher
12 Years Excellent
Recommended additional case study templates.
Provide templates for various business scenarios.
P7
Empirical Researcher
7 Years Not Satisfied
Provided result was not expected.
Need to improve the tool scalability.
P8
Theoretical Researcher
9 Years Very Good
Requested support for more languages.
Add additional language capabilities.
P9
SE Researcher
4 Years Excellent
Impressed with multi-method approach.
Continue to refine multi-method capabilities.
P10
Empirical Researcher
15 Years Fair
Suggested a feature for better paper selection.
Create open source model for data extractions.
@@ 5 Discussions
The results derived from the implementation of our multi-AI agent model for SLR have been both encouraging and insightful. The model successfully automated key components of the SLR process, including the generation of search strings, the selection and filtering of relevant literature, and the summarization of key findings. This automation significantly reduced the time and effort typically required in conducting SLRs, while maintaining, and in some aspects enhancing, the accuracy and comprehensiveness of the review. The model’s ability to process and analyze large volumes of text rapidly, and its precision in identifying relevant studies, demonstrated the substantial potential of integrating LLMs in academic research.
@@
The implications of these results are far-reaching. Firstly, the model presents a valuable tool for researchers across various fields, reducing the barriers to conducting comprehensive literature reviews. This efficiency can accelerate the pace of research and discovery, enabling scholars to focus on more complex and creative aspects of their work. Additionally, the model’s standardization of the SLR process can potentially lead to more consistent and replicable research outcomes, a cornerstone in scientific research. The reduction in manual effort also opens up opportunities for researchers with limited resources or those facing constraints such as time pressures, broadening the scope of who can conduct thorough literature reviews.
Looking ahead, the future impact of our model is poised to be significant. A key milestone will be our attendance at the Sanner Conference on March 12th, where we will showcase our developed tool to a diverse audience. This event will not only serve as a platform to demonstrate the capabilities of our model but also as a critical opportunity to gather feedback from a wide range of users. This feedback will be invaluable in refining and enhancing the model further. Understanding how the model performs in real-world scenarios and gathering diverse perspectives will enable us to tailor it more closely to the needs of the research community. Following this, we plan to implement updates and improvements based on this feedback, ensuring that our tool remains at the forefront of innovation in SLR automation. The continued development and adaptation of our model in response to user input will ensure its relevance and utility in the ever-evolving landscape of academic research.
@@
In addition, the long-term impact of our work on researchers and the broader academic community is expected to be substantial. Our model represents a paradigm shift in how SLRs are conducted, offering a tool that is not only efficient but also adaptable to the evolving needs of researchers. One significant future impact is the democratization of research. By simplifying the SLR process, our tool makes high-quality literature reviews accessible to a wider range of researchers, including those from institutions with fewer resources or those new to the field. This accessibility could lead to a more diverse range of voices and perspectives in academic research, enriching the field as a whole.
@@
Furthermore, the model’s efficiency in handling large volumes of data makes it an invaluable asset in fields where literature is vast and rapidly growing, such as biomedical research, environmental studies, and technology. Researchers in these fields can stay abreast of the latest developments more effectively, ensuring their work is informed by the most current and comprehensive data available.
@@
In the domain of interdisciplinary research, our model can facilitate the synthesis of information across different fields, potentially leading to novel insights and innovations. By efficiently collating and analyzing diverse sets of literature, the tool can help uncover connections between disciplines that might otherwise be overlooked.
@@
The long-term adaptation of our model based on user feedback and technological advancements will also ensure its ongoing relevance. Continuous updates will allow the model to incorporate the latest AI advancements, further enhancing its capabilities and ensuring it remains a cutting-edge tool for SLR. Moreover, the model’s potential for customization will allow it to cater to the specific needs of different research domains. This customized approach means that the tool can be fine-tuned to deliver more targeted and relevant results, depending on the specific requirements of the research question or field.

### v2

@@ Abstract
Systematic literature review (SLR) is foundational to evidence-based research, enabling scholars to identify, classify, and synthesize existing studies to address specific research questions. Conducting an SLR is, however, largely a manual process. In recent years, researchers have made significant progress in automating portions of the SLR pipeline to reduce the effort and time required for high-quality reviews; nevertheless, there remains a lack of AI-agent-based systems that automate the entire SLR workflow. To this end, we introduce a novel multi-AI-agent system designed to fully automate SLRs. Leveraging large language models (LLMs), our system streamlines the review process to enhance efficiency and accuracy. Through a user-friendly interface, researchers specify a topic; the system then generates a search string to retrieve relevant academic papers. Next, an inclusion/exclusion filtering step is applied to titles relevant to the research area. The system subsequently summarizes paper abstracts and retains only those directly related to the field of study. In the final phase, it conducts a thorough analysis of the selected papers with respect to predefined research questions. This paper presents the system, describes its operational framework, and demonstrates how it substantially reduces the time and effort traditionally required for SLRs while maintaining comprehensiveness and precision. The code for this project is available at: https://github.com/GPT-Laboratory/SLR-automation .
@@ 1 Introduction
In this context, our proposed system utilizes the capabilities of LLMs to automate the whole SLR process. We developed a LLM-based multi-agent system that automates each step of the SLR, from the initial literature search to the final analysis. The system begins with a simple user input–researchers enter their topic into a designated text box. This input is then processed by the LLM, which generates a precise search string tailored to retrieve the most relevant academic papers. The system’s next phase involves an intelligent filtering mechanism. It applies an inclusive and exclusive theory, screening titles and abstracts to retain only those studies that are directly relevant to the specified research area.
The final stage of our system autonomously summarizes the abstracts of the selected papers, ensuring that only content pertinent to the research questions is retained. It introduces a level of precision and consistency in data analysis that is challenging to achieve manually. Finally, the proposed system conducts an in-depth analysis of the selected papers, aligning its findings directly with the research questions. This comprehensive approach ensures that the final output is not only a reflection of the vast array of literature available but also a focused and relevant resource tailored to the specific needs of the researcher. Our system, therefore, stands as a testament to the potential of integrating advanced AI technologies in academic research methodologies.
@@ 2 Related Work
Despite these advancements in automating the SLR process, there remains a notable gap in the complete automation of SLRs using LLMs. Addressing this gap, we have developed a novel approach: a multi-agent system based on LLMs. This innovative system is designed to fully automate the SLR process, utilizing the advanced capabilities of LLMs to efficiently manage and synthesize vast amounts of data, which is a significant step forward in the field of automated literature reviews.
@@ 3 Research Method
This research aims to investigate how an LLM-based multi-agent system can be utilized to automate the entire process of SLRs. Below, we discuss how our LLM-based multi-agent system collaborates and performs such tasks. We have formulated a research questions (RQ):
RQ. How does a LLM-based multi-agent system transform traditional methodologies to automate the systematic literature review process in SE?
Motivation: The motivation for the research question arises from the need to enhance the efficiency and effectiveness of literature review processes in the rapidly evolving field of SE. Traditional methods of conducting literature reviews are often time-consuming and labor-intensive, potentially leading to delays in research progress and the dissemination of new knowledge. The integration of LLMs-based agents promises a paradigm shift, potentially automating and streamlining these processes. By exploring the transformation brought about by an LLM-based multi-agent system, this research seeks to reduce the time and effort required for comprehensive literature reviews and also to increase the accuracy and scope of these reviews. This could result in more timely and informed research outcomes in SE, a field where staying abreast of current trends, methodologies, and discoveries are crucial for technological advancement and innovation.
3.1 LLM-based Agent Assisted Systematic Literature Review
The section focuses on research methodology for developing an LLM-based multi-agent system. This system is specifically engineered to automate the entire process of conducting SLRs. The innovation lies in its ability to transform a given research topic into a comprehensive review through a series of automated, interconnected steps. Each step is managed by a specialized agent within the system, working collaboratively to ensure a seamless and efficient literature review process. In Figure LABEL:fig:_multi-agent_system, we illustrate how agents collaborate with each other to generate a response. Below, we also provide the detailed functionality of each agent within this multi-agent system.
@@ 3.1.1 Planner agent
The first agent is dedicated to generating research questions and purpose and search strings. Upon receiving a research objective from the end-user, For the search string we provide research questions and objective to the agent, it generates a search string. this agent employs advanced language understanding algorithms to interpret the topic’s key elements. The system, designed for deep semantic understanding, analyzes the topic to extract key concepts, themes, and terminologies. It then utilizes its extensive training on diverse textual data to construct a precise and comprehensive search string. This string is formulated by combining relevant keywords, synonyms, and technical terms that capture the essence of the research question. Furthermore, the algorithm is adept at understanding context and varying semantic structures, enabling it to refine the search string to match specific research domains. The generated search string is crucial in accurately retrieving relevant literature from various academic databases. By ensuring that the initial search is both thorough and focused, the agent significantly enhances the efficiency and quality of the literature collection process. This sets a solid foundation for the subsequent stages of the SLR, where the depth and breadth of the collected literature play a crucial role.
@@ 3.1.3 Data extraction agent
The third agent in our system is tasked with refining the literature using inclusion and exclusion criteria based on the research objectives. Initially, it employs our LLM algorithm to analyze the titles of retrieved papers, discerning their relevance to the research topic. This step involves text analysis, where the LLM algorithm identifies key terms and concepts that align with the research objectives. By applying these predefined rules, the agent effectively filters out irrelevant material, ensuring the literature review remains focused and pertinent to the research questions.
@@ 3.1.4 Data compilation agent
The final agent in our multi-agent system is responsible for analyzing the synthesized data in relation to the research questions and objectives. It assesses trends, identifies gaps in the literature, and draws conclusions based on the aggregated information. This agent also prepares a report that summarizes the findings of the literature review, providing a clear and concise overview of the research landscape for the given topic.
Each agent in the LLM-based multi-agent system plays a vital role in automating the systematic literature review process. From generating search strings to reporting findings, the agents work in a coordinated manner to ensure a thorough, efficient, and accurate review. This methodology represents a significant advancement in the way SLRs are conducted, offering a more streamlined and effective approach to academic research.
@@ 4 Results
This section presents the results obtained from implementing an AI-agent-based system aimed at automating the SLR process in SE. The findings are provided below in detail.
4.1 LLM Based Multi-Agent System (RQ)
Our research introduces an LLM-based multi-agent system that redefines the conventional approach to SLRs in SE. The multi-agent system developed for automating the SLR process has demonstrated its efficacy through a structured and sequential workflow. The process begins with the input of a research topic, as depicted in Figure 2.
@@ 4.1 LLM Based Multi-Agent System (RQ)
Upon receiving the topic, the system systematically generates a pertinent set of research questions. As illustrated in Figure 2, it formulated questions like ’How have large language models been utilized in various aspects of the software development process?’ and ’What challenges and limitations exist in the adoption and implementation of large language models in software development?’ These questions play a crucial role in guiding the literature search and analysis process. Subsequent to the research question formulation, the system proceeds to generate a search string. In this case, the search string “large language models OR software development” was created, coupled with a specified year to narrow down the search scope, thereby enhancing the relevance and precision of the search results.
@@
The subsequent phase involves retrieving papers that match the generated search string. This feature of the system is specifically designed to fetch research papers from various databases, as demonstrated by the list of papers shown in Figure 3. For this demonstration, we focused solely on papers published in the year 2023, setting the system to retrieve only 10 papers from that year, all relevant to this field. The tool efficiently compiles relevant information such as the title, author, publication URL, journal name, DOI, paper type, affiliation country, and the affiliation institution. Moreover, the system is equipped with the capability to apply inclusive and exclusive criteria based on titles, which further refines the search results to ensure only the most pertinent literature is considered for review. As shown in Figure 4, only three papers were selected for an in-depth analysis.
Finally, the system extracts data based on the formulated RQs. This advanced feature is exemplified in the demo where detailed answers are provided for the previously generated RQs. For instance, the answer to the first RQs discusses the varied applications of large language models in the software development lifecycle and highlights specific instances of their use, like the inference from the paper “InferLink End-to-End Program Repair with Large Language Models.”
@@
In conclusion, the proposed system has showcased its ability to streamline the laborious process of literature review, from defining the scope of the research to extracting and synthesizing data pertinent to the research questions. The demonstration affirms the system’s potential in significantly reducing the time and effort conventionally required for conducting systematic literature reviews.
@@ 5 Discussions
The results derived from the implementation of our multi-AI agent system for SLR have been both encouraging and insightful. The system successfully automated key components of the SLR process, including the generation of search strings, the selection and filtering of relevant literature, and the summarization of key findings. This automation significantly reduced the time and effort typically required in conducting SLRs, while maintaining, and in some aspects enhancing, the accuracy and comprehensiveness of the review. The system’s ability to process and analyze large volumes of text rapidly, and its precision in identifying relevant studies, demonstrated the substantial potential of integrating LLMs in academic research.
@@
The implications of these results are far-reaching. Firstly, the system presents a valuable tool for researchers across various fields, reducing the barriers to conducting comprehensive literature reviews. This efficiency can accelerate the pace of research and discovery, enabling scholars to focus on more complex and creative aspects of their work. Additionally, the system’s standardization of the SLR process can potentially lead to more consistent and replicable research outcomes, a cornerstone in scientific research. The reduction in manual effort also opens up opportunities for researchers with limited resources or those facing constraints such as time pressures, broadening the scope of who can conduct thorough literature reviews.
@@
In addition, the long-term impact of our work on researchers and the broader academic community is expected to be substantial. Our system represents a paradigm shift in how SLRs are conducted, offering a tool that is not only efficient but also adaptable to the evolving needs of researchers. One significant future impact is the democratization of research. By simplifying the SLR process, our tool makes high-quality literature reviews accessible to a wider range of researchers, including those from institutions with fewer resources or those new to the field. This accessibility could lead to a more diverse range of voices and perspectives in academic research, enriching the field as a whole.
@@
Furthermore, the system’s efficiency in handling large volumes of data makes it an invaluable asset in fields where literature is vast and rapidly growing, such as biomedical research, environmental studies, and technology. Researchers in these fields can stay abreast of the latest developments more effectively, ensuring their work is informed by the most current and comprehensive data available.
@@
In the domain of interdisciplinary research, our system can facilitate the synthesis of information across different fields, potentially leading to novel insights and innovations. By efficiently collating and analyzing diverse sets of literature, the tool can help uncover connections between disciplines that might otherwise be overlooked.
@@
The long-term adaptation of our system based on user feedback and technological advancements will also ensure its ongoing relevance. Continuous updates will allow the system to incorporate the latest AI advancements, further enhancing its capabilities and ensuring it remains a cutting-edge tool for SLR. Moreover, the system’s potential for customization will allow it to cater to the specific needs of different research domains. This customized approach means that the tool can be fine-tuned to deliver more targeted and relevant results, depending on the specific requirements of the research question or field.